OpenStudy (anonymous):

-7/20f+2/5=1/4f what does f equal (show work) 4

myininaya (myininaya):

\[-\frac{7}{20}f+\frac{2}{5}=\frac{1}{4}f\] get your f's together add 7f/20 on both sides \[\frac{2}{5}=\frac{1}{4}f+\frac{7}{20}f\] find a common denominator \[\frac{2}{5}=\frac{5}{20}f+\frac{7}{20}f\] \[\frac{2}{5}=\frac{12}{20}f\] we can simplify 12/12 \[\frac{2}{5}=\frac{3}{5}f\] \[\frac{5}{3}*\frac{2}{5}=f\]

OpenStudy (anonymous):

2/5=1/4f+7/20f 2/5=5/20f+7/20f 2/5=12/20f 4f=60 f=15

myininaya (myininaya):

i get f=2/3 lets check

OpenStudy (anonymous):

40f=60 f=3/2

myininaya (myininaya):

a7med did you do f=20/12*2/5

OpenStudy (anonymous):

20f*2=5*12

myininaya (myininaya):

so you tried to cross multipliy so it should be 5*12f=2*20

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

no, u 20f*2=5*12 i'm sure

myininaya (myininaya):

we have 12f/20=2/5 to get f by itself we multiplied by the reciprocal of 12/20 on both sides giving us (20/12)(12f/20)=(20/12)(2/5) f=40/60=2/3

myininaya (myininaya):

we can also check f=2/3 so we plug 2/3 whereever we see an f in the orignal problem (-7/20)(2/3)+2/5=(1/4)(2/3) -14/60+2/5=2/12 -7/30+2/5=2/12 -7/30+12/30=2/12 5/30=2/12 1/6=1/6 2/3 is definitely correct!